# SeqAfford: Sequential 3D Affordance Reasoning via Multimodal Large Language Model

### v1

@@ Abstract
3D affordance segmentation aims to link human instructions to touchable regions of 3D objects for embodied manipulations. Existing efforts typically adhere to single-object, single-affordance paradigms, where each affordance type or explicit instruction strictly corresponds to a specific affordance region and are unable to handle long-horizon tasks. Such a paradigm cannot actively reason about complex user intentions that often imply sequential affordances. In this paper, we introduce the Sequential 3D Affordance Reasoning task, which extends the traditional paradigm by reasoning from cumbersome user intentions and then decomposing them into a series of segmentation maps. Toward this, we construct the first instruction-based affordance segmentation benchmark that includes reasoning over both single and sequential affordances, comprising 180K instruction-point cloud pairs. Based on the benchmark, we propose our model, SeqAfford, to unlock the 3D multi-modal large language model with additional affordance segmentation abilities, which ensures reasoning with world knowledge and fine-grained affordance grounding in a cohesive framework. We further introduce a multi-granular language-point integration module to endow 3D dense prediction. Extensive experimental evaluations show that our model excels over well-established methods and exhibits open-world generalization with sequential reasoning abilities.

### v2

@@ Abstract
3D affordance segmentation aims to link human instructions to touchable regions of 3D objects for embodied manipulations. Existing efforts typically adhere to single-object, single-affordance paradigms, where each affordance type or explicit instruction strictly corresponds to a specific affordance region and are unable to handle long-horizon tasks. Such a paradigm cannot actively reason about complex user intentions that often imply sequential affordances. In this paper, we introduce the Sequential 3D Affordance Reasoning task, which extends the traditional paradigm by reasoning from cumbersome user intentions and then decomposing them into a series of segmentation maps. Toward this, we construct the first instruction-based affordance segmentation benchmark that includes reasoning over both single and sequential affordances, comprising 180K instruction-point cloud pairs. Based on the benchmark, we propose our model, SeqAfford, to unlock the 3D multi-modal large language model with additional affordance segmentation abilities, which ensures reasoning with world knowledge and fine-grained affordance grounding in a cohesive framework. We further introduce a multi-granular language-point integration module to endow 3D dense prediction. Extensive experimental evaluations show that our model excels over well-established methods and exhibits open-world generalization with sequential reasoning abilities. Code and Dateset are available at: https://github.com/hq-King/SeqAfford.
@@ 6 Conclusion
Supplementary Materials
[…]
Supplementary organization:
Appendix A Code and Dataset
We will release our code and dataset once the paper is accepted.
Appendix B Textual Prompt Format
This section outlines the specific design of the Textual Prompt component included in the four generative paradigms demonstrated earlier. By designing the Textual Prompt, it helps to enhance GPT-4’s understanding and generalization ability for the relevant tasks, providing suitable introductions for subsequent tasks. We have utilized the principles of prompt engineering to design modules such as Role, Task, Example, and Instruction for the Textual prompt. Additionally, we have formatted the input in JSON format to improve GPT-4’s comprehension of the text, standardizing the input of Object Name and Affordance Type.We show the details in Fig. A1.
B.1 Role
In the role module, we have preset the textual Prompt template as follows: "Role: You are an analytical assistant specializing in robotic affordance grounding. Your expertise is in creating questions that facilitate the training of robotic affordance grounding, enabling robots to reason about task execution, such as determining the appropriate part of an object to grasp."
B.2 Task
In the task module, we have preset the textual Prompt template as follows: "Task: Description: You will be provided with the name of an object. The robot is expected to use this tool to perform a variety of everyday tasks. Along with the tool name, you will receive an affordance type that the object can afford and a list of questions that have already been generated for this tool. Guidelines: Diversity: Aim for a wide range of tasks, ensuring that there is no overlap with previous ones. Daily Tasks: questions should be common and representative of the ones encountered in daily life. Leakage Avoidance: Ensure that the generated questions tasks can only afford the given affordance."
B.3 Examples
Examples illustrate the types of questions that can be generated based on the object name and affordance type. These examples guide the creator of the prompt in formulating new and diverse questions. In the task module, we have preset the textual Prompt template as follows: "Examples: When cutting a rope with these scissors, which part of it should your palm touch? If you want to lift the bag, at which point is your finger most likely to carry it? Point out the areas on the microwave ideal for opening. How would you grasp the hat to best maintain its condition? If you want to boil water, at which points on the tap would you open the water valve? And so on…".
B.4 Instruction
In the task module, we have preset the textual Prompt template as follows: "Instruction: With the provided OBJECT NAME: ’+object+’ and AFFORDANCE TYPE: ’+affordance types+’, generate fifteen new affordance grounding question tasks. Use the HISTORY of generated tasks as a reference to ensure compliance with the diversity guideline. Output should be in the JSON format."
B.5 Template
Here is the template that encapsulates all the elements of the textual prompt:
Figure A1: textual Prompt Format
B.6 Genereted Instructions Example
We illustrated detailed examples of our generated instructions in Fig. A2, A3, A4, A5
Figure A2: Genereted Instructions Example:Textual Prompt
Figure A3: Genereted Instructions Example:Textual Prompt with Mesh Images
Figure A4: Genereted Instructions Example:Textual Prompt with Mesh and HOI Images
Figure A5: Genereted Instructions Example:Textual Prompt with Mesh Images and Scenarios Descriptions
Appendix C Format of Questions and and Answers
We have several formats of input, we show the Question List template , Explanatory Question List template and Answer List template in Fig. A6
Figure A6: The details of our templates
Table A1: Dataset Statistics. This table showcases the distribution of our instruction-tuning benchmark for the Sequential Affordance Reasoning task, which introduces a wide range of variances and realistic simulations, including seen/unseen settings, single/sequential scenarios.
Task Setting Train Test
Shapes Instruction-Point Cloud Pairs Shapes Instruction-Point Cloud Pairs
Single Seen 16084 157,820 2287 4566
Single Unseen 14307 124,140 2287 4566
Sequential Seen 8156 13393 2,946 7454
Appendix D Dataset Details
Our dataset comprises 162,386 instruction-point cloud pairs in the single affordance segmentation setting and 20,847 pairs in the sequential affordance segmentation setting, making up a total of 18,371 point cloud instances across 23 object categories. We have visualized word clouds for the object categories, affordance types, and the introduction in our dataset, demonstrating the richness of our dataset. Fig. A7 shows the distribution of the generated question length and Fig. A8 shows the words cloud of the generated instructions.
Figure A7: Question Length Distribution
(a) Instructions
(b) Affordance Types
(c) Object Categories
Figure A8: Word clouds of (a) instructions, (b) affordance types, and (c) object categories.
Figure A9: Upsample. We show the details of the Upsample method.Specifically, we use FPS and feature propagation to get the hierarchical point cloud features and upsample the sparse features into dense features
Appendix E Detailed Methodology
E.1 Point Upsample
Fig. A9 shows the detailed method of our upsample ways. Here, we adopt Uni3D[41] as our 3D encoder, which is capable of encoding point clouds to obtain semantically rich point cloud features $Fp_{sparse}$. To perform dense prediction tasks, we need to obtain finer-grained point cloud features from the sparse point cloud features $Fp_{sparse}$. Following PointNet++[26] and DGCNN[34], we utilized their feature propagation techniques to build our feature propagation process. Specifically, we extract features from the intermediate layers of the 3D encoder and use Farthest Point Sampling (FPS) to sample different numbers of points from the point cloud:
$P2=FPS(X_{point}),$ (5)
$P3=FPS(X_{point}).$ (6)
Feature propagation is then applied to obtain features $f1$ and $f2$:
$f1=FP(P1,P2,H8),$ (7)
$f2=FP(P1,P3,H4).$ (8)
Next, we follow DGCNN[34] using its upsample technique to obtain $f1^{{}^{\prime}}$ and $f2^{{}^{\prime}}$:
$f1^{{}^{\prime}}=Up(P2,f1,P2.H8),$ (9)
$f2^{{}^{\prime}}=Up(P3,f2,P2,f1^{{}^{\prime}})$ (10)
E.2 Mutil-Object
Here, we describe the detailed implementation of the Mutil-Object affordance reasoning. During training, we add $<SEG>$ token after the object name, and our MLLM will learn the paragrim and when it needs to generate text output, it will also follow this paragtim. We extract the nouns from the output of the language model, we then use the order of these nouns as the order of the model affordance reasoning. This is justified because the large language model is able to output the correct text sequence.
Appendix F More Visualization Results
Here, we show more qualitative results of our model. Fig. A10 shows the additional visualization results of Single Affordance Reasoging, and Fig. A11 shows the additional visualization results of
Figure A10: More Qualitative Visualization Results of Sequential Affordance Reasoning
Figure A11: More Qualitative Visualization Results of Sequential Affordance Reasoning
Appendix G Comparison with the Previous SOTA Method
We show the qualitative comparison with LASO[12]: obviously, when the question is complex, LASO[12] failed to understand the intention and do not perform well in connecting the affordance with the intention.
Figure A12: Comparison with the Previous SOTA Method